OpenStudy (anonymous):

solve equation. -4= 7over33 x x=132 over 7 x =-132over 7 x=-7over 132 x=-33over28 which equation is and identity 8-(5x+2)=-5x-6 b. 7z+10-z=8z-2(z-5) c. 8m-4=5m=8-m d. 6y+5+6y-5

OpenStudy (anonymous):

-4=7/33x so we start by multiplying 33x on both sides -4(33x)=7 then distribute -132x=7 divide by -132 x=-7/132 C

OpenStudy (anonymous):

and can you clarify the 2nd question please

OpenStudy (anonymous):

which one is an identity*

OpenStudy (anonymous):

ok hold on a bit

OpenStudy (anonymous):

an identity means the when the equations on both sides but into their final form, they are identical so lets look at each answer a. 8-(5x+2)=-5x-6 8-5x-2=-5x-6 -5x+6=-5x-6 so no b. 7z+10-z=8z-2(z-5) 6z+10=8z-2z+10 6z+10=6z+10 yes but lets check the others c. 8m-4=5m=8-m uhhh? d. 6y+5+6y-5 this cancels each other and equals 0 so no

OpenStudy (anonymous):

so B